# Fairness-Aware Domain Generalization under Covariate and Dependence Shifts

### v1

@@ 5 Experiments
FDDG 0.87$\pm$0.09 / 0.53$\pm$0.01 / 62.56$\pm$2.25 0.94$\pm$0.05 / 0.52$\pm$0.01 / 93.36$\pm$1.70 0.93$\pm$0.03 / 0.53$\pm$0.02 / 93.43$\pm$0.73 0.92 / 0.53 / 83.12
Table 4: Ablation studies results on FairFace.
DP $\uparrow$ / AUC $\downarrow$ / Accuracy $\uparrow$
Methods (B, 0.91) (L, 0.48) (W, 0.49) Avg
FDDG w/o sf 0.68$\pm$0.18 / 0.57$\pm$0.02 / 93.07$\pm$0.68 0.47$\pm$0.07 / 0.63$\pm$0.01 / 92.62$\pm$0.93 0.35$\pm$0.26 / 0.58$\pm$0.01 / 92.18$\pm$0.46 0.49 / 0.59 / 93.08
FDDG w/o $T$ 0.83$\pm$0.08 / 0.56$\pm$0.01 / 92.81$\pm$0.81 0.53$\pm$0.03 / 0.59$\pm$0.01 / 91.19$\pm$0.57 0.52$\pm$0.23 / 0.59$\pm$0.01 / 90.78$\pm$0.31 0.58 / 0.59 / 91.65
FDDG w/o fc 0.59$\pm$0.16 / 0.58$\pm$0.01 / 92.92$\pm$1.35 0.40$\pm$0.07 / 0.70$\pm$0.02 / 92.96$\pm$0.85 0.34$\pm$0.08 / 0.72$\pm$0.03 / 91.88$\pm$0.67 0.42 / 0.66 / 93.01
@@ 6 Conclusion
Differing from established domain generalization research, which attributes distribution shifts from source domains to target domains to concept shift and covariate shift, we introduce a novel form of distribution shift known as dependence shift. This shift pertains to varying dependence patterns between model outcomes and sensitive attributes in different domains. Furthermore, we introduce a novel approach to fairness-aware learning designed to tackle the challenges of domain generalization when confronted with both covariate shift and dependence shift simultaneously. In our pursuit of learning a fairness-aware invariant classifier across domains, we assert the existence of an underlying transformation model that can transform instances from one domain to another. This model plays a crucial role in achieving fairness-aware domain generalization by generating instances in synthetic domains characterized by novel data styles and fair dependence patterns. We present a practical and tractable algorithm, accompanied by comprehensive theoretical analyses and exhaustive empirical studies. We showcase the algorithm’s effectiveness through rigorous comparisons with state-of-the-art baselines.
@@
Appendix A Related Works
Domain generalization. Addressing the challenge of domain shift and the absence of OOD data has led to the introduction of several state-of-the-art methods in the domain generalization field [24, 8, 13, 12]. These methods are designed to enable deep learning models to possess intrinsic generalizability, allowing them to adapt effectively from one or multiple source domains to target domains characterized by unknown distributions [35]. They encompass various techniques, such as aligning source domain distributions to facilitate domain-invariant representation learning [29], subjecting the model to domain shift during training through meta-learning [27], and augmenting data with domain analysis, among others [36], and so on. In the context of the number of source domains, a significant portion of research [13, 12, 37] has focused on the multi-source setting. This setting assumes the availability of multiple distinct but relevant domains for the generalization task. As mentioned in [37], the primary motivation for studying domain generalization is to harness data from multiple sources in order to unveil stable patterns. This entails learning representations that are invariant to the marginal distributions of data features, all while lacking access to the target data. Nevertheless, existing domain generalization methods tend to overlook the aspect of learning with fairness, where group fairness dependence patterns may not undergo changes across domains.
Fairness learning for changing environments. Two primary research directions aim to tackle fairness-aware machine learning in dynamic or changing environments. The first approach involves equality-aware monitoring methods [38, 39, 6, 40, 41, 42, 43], which strive to identify and mitigate unfairness in a model’s behavior by continuously monitoring its predictions. These methods adapt the model’s parameters or structure when unfairness is detected. However, a significant limitation of such approaches is their assumption of invariant fairness levels across domains, which may not hold in real-world applications. The second approach [5, 4] focuses on evaluating a model’s fairness in a dynamic environment by treating shifted fairness levels as domain labels. However, it does not take into account distribution shifts in non-sensitive features.
In response to these limitations, this paper adopts a novel approach by attributing the distribution shift from source to target domains to both covariate shift and fairness dependence shift simultaneously. We aim to train a fairness-aware invariant classifier that can generalize effectively across domains, ensuring robust performance in terms of both model accuracy and maintaining fair dependence between predicted outcomes and sensitive attributes even under these shifts.
Appendix B Notations
@@
$\mathcal{C}$ parameterized latent space for content factors
$\mathcal{S}$ parameterized latent space for style factors
$\mathcal{A}$ parameterized latent space for sensitive factors
$\mathbf{c}$ content factor
@@
$d[\cdot]$ distance metric on output space
$\mathcal{D}$ data batch
@@
$\mathcal{F}$ model space of classifier
@@
$g(Y,Z)$ fairness metric on random variables $Y$ and $Z$
@@
$e$ data domain
$\mathcal{E}$ set of data domains
@@
$\delta,\epsilon$ expectation of the relaxed constraint
$h$ sensitive label classifier
$\hat{z}$ sensitive label predicted by $h$
@@
Appendix C Experimental Settings
C.1 Datasets.
We consider four datasets: ccMNIST, FairFace, YFCC100M-FDG, and New York Stop-and-Frisk (NYSF) to evaluate our FDDG against state-of-the-art baseline methods, where NYSF is a tabular data and the other three are image datasets.
(a) ccMNIST is a domain generalization benchmark created by colorizing digits and the backgrounds of the MNIST dataset [44]. ccMNIST consists of images of handwritten digits from 0 to 9. Similar to ColoredMNIST [8], for binary classification, digits are labeled with 0 and 1 for digits from 0-4 and 5-9, respectively. ccMNIST contains three data domains, each characterized by a different digit color (i.e., red, green, blue) with 70,000 images. Each image has a black or white background color as the sensitive label. The domains are constructed so that each domain has a different correlation between the class label and sensitive attribute (digit background colors), specifically $0.9$ for the red domain, $0.7$ for the green domain, and 00 for the blue domain.
(b) FairFace [21] is a dataset that contains a balanced representation of different racial groups. It includes 108,501 images from seven racial categories: Black (B), East Asian (E), Indian (I), Latino (L), Middle Eastern (M), Southeast Asian (S), and White (W). In our experiments, we set each racial group as a domain, gender as the sensitive label, and age ($\geq$ or $<$ 50) as the class label.
(c) YFCC100M-FDG is an image dataset created by Yahoo Labs and released to the public in 2014. It is randomly selected from the YFCC100M [22] dataset with a total of 90,000 images. For domain variations, YFCC100M-FDG is divided into three domains. Each contains 30,000 images from different year ranges, before 1999 ($d_{0}$), 2000 to 2009 ($d_{1}$), and 2010 to 2014 ($d_{2}$). The outdoor or Indoor tag is used as the binary class label for each image. Latitude and longitude coordinates, representing where images were taken, are translated into different continents. The continent North-America or non-North-America is used as the sensitive label (related to spatial disparity).
(d) NYSF [23] is a real-world dataset on policing in New York City in 2011. It documents whether a pedestrian who was stopped on suspicion of weapon possession would in fact possess a weapon. NYSF consists of records collected in five different sub-cities, Manhattan (M), Brooklyn (B), Queens (Q), Bronx (R), and Staten (S). We use cities as different domains. This data had a pronounced racial bias against African Americans, so we consider race (black or non-black) as the sensitive attribute.
C.2 Baselines.
We compare the performance of our FDDG with 17 baseline methods that fall into three main categories:
  • 12 state-of-the-art domain generalizations methods (RandAug, ERM [24], IRM [8], GDRO [25], Mixup [26], MLDG [27], CORAL [28], MMD [29], DANN [30], CDANN [31], DDG [13], and MBDG [12]);
  • 3 state-of-the-art fairness-aware learning methods in changing environments (EIIL [4], FarconVAE [5], and FATDM [6]);
  • 2 naive fairness-aware variants of DDG and MBDG, named DDG-FC and MBDG-FC, respectively, by simply adding fairness constraints in Defn. 1 to their classifiers.
Notice that the settings of EIIL and FarconVAE are different from this paper. Both methods characterize domain shift by a different level of correlation between the class label and sensitive features but completely ignore the variation in data features.
C.3 Evaluation Metrics.
Three metrics are used for evaluation, and two of them are for fairness quantification.
@@ Appendix C Experimental Settings
    $\text{DP}=k,\>\text{if DP}\leq 1;\text{DP}=1/k,\>\text{otherwise},\quad\text{where}\>k=\mathbb{P}(\hat{Y}=1|Z=-1)/\mathbb{P}(\hat{Y}=1|Z=1)$
    This is also known as a lack of disparate impact [45]. A value closer to 1 indicates fairness.
  • The Area Under the ROC Curve (AUC) [34] varies from zero to one, and it is symmetric around 0.5, which represents random predictability or zero bias effect on predictions.
    $\text{AUC}=\frac{\sum_{(\mathbf{x}_{i},z=-1,y_{i})\in\mathcal{D}_{-1}}\sum_{(\mathbf{x}_{j},z=1,y_{j})\in\mathcal{D}_{1}}I\big{(}\mathbb{P}(\hat{y}_{i}=1)>\mathbb{P}(\hat{y}_{j}=1)\big{)}}{|\mathcal{D}_{-1}|\times|\mathcal{D}_{1}|}$
@@
C.4 Learning the Transformation Model
One goal of the transformation model $T$ is to disentangle an input instance from source domains into three factors in latent spaces by learning a set of encoders $E=\{E^{m},E^{s},E^{c},E^{a}\}$ and decoders $G=\{G^{i},G^{o}\}$ parameterized by $\{\bm{\theta}_{m},\bm{\theta}_{s},\bm{\theta}_{c},\bm{\theta}_{a}\}\in\Theta$ and $\{\bm{\phi}_{i},\bm{\phi}_{o}\}\in\Phi$, respectively. As shown in Fig. 5, the learning process of $T$ consists of two levels, an outer level and an inner level, where each level is associated with an auto-encoder system factorizing its corresponding input into two factors within two separated latent spaces. Specifically, in the outer level, an instance is first encoded to a semantic factor $\mathbf{m}\in\mathcal{M}$ and a style factor $\mathbf{s}\in\mathcal{S}$ through the corresponding encoders $E^{m}:\mathcal{X}\times\Theta\rightarrow\mathcal{M}$ and $E^{s}:\mathcal{X}\times\Theta\rightarrow\mathcal{S}$, respectively. In the inner level, the semantic factor $\mathbf{m}$ is further encoded to a content factor $\mathbf{c}\in\mathcal{C}$ and a sensitive factor $\mathbf{a}\in\mathcal{A}$, through encoders $E^{c}:\mathcal{M}\times\Theta\rightarrow\mathcal{C}$ and $E^{a}:\mathcal{M}\times\Theta\rightarrow\mathcal{A}$. For data reconstruction, two decoders $G^{i}:\mathcal{C}\times\mathcal{A}\times\Phi\rightarrow\mathcal{M}$ and $G^{o}:\mathcal{M}\times\mathcal{S}\times\Phi\rightarrow\mathcal{X}$ are introduced in the inner and outer levels. Inspired by image-to-image translation in computer vision [18, 19], our total loss function of learning such encoders and decoders comprises three components: a bidirectional reconstruction loss, a sensitive label prediction loss, and an adversarial loss.
[…]
$\mathcal{L}_{recon}^{data}=\underbrace{\mathbb{E}_{\mathbf{x}\sim p(\mathbf{x})}\big{[}\big{\lVert}G^{o}\big{(}\hat{\mathbf{m}},E^{s}(\mathbf{x})\big{)}-\mathbf{x}\big{\rVert}_{1}\big{]}}_{\text{data reconstruction}}+\underbrace{\mathbb{E}_{\mathbf{m}\sim p(\mathbf{m})}\big{[}\big{\lVert}G^{i}\big{(}E^{c}(\mathbf{m}),E^{a}(\mathbf{m})\big{)}-\mathbf{m}\big{\rVert}_{1}\big{]}}_{\text{reconstruction of semantic factors $\mathbf{m}$ (inner level)}}$
where $\hat{\mathbf{m}}=G^{i}(\mathbf{c,a})=G^{i}\big{(}E^{c}(E^{m}(\mathbf{x})),E^{a}(E^{m}(\mathbf{x}))\big{)}$; $p(\mathbf{m})$ is given by $\mathbf{m}=E^{m}(\mathbf{x})$ and $\mathbf{x}\sim p(\mathbf{x})$. For factor reconstruction, $\mathbf{m,s,c}$ and $\mathbf{a}$ are encouraged to be reconstructed through some latent factors randomly sampled from the prior distributions.
$\mathcal{L}_{recon}^{factor}=\underbrace{\mathbb{E}_{\mathbf{c}\sim p(\mathbf{c}),\mathbf{a}\sim\mathcal{N}(0,\mathbf{I}_{a})}\big{[}\big{\lVert}E^{c}\big{(}G^{i}(\mathbf{c,a})\big{)}-\mathbf{c}\big{\rVert}_{1}\big{]}}_{\text{reconstruction of content factors $\mathbf{c}$}}+\underbrace{\mathbb{E}_{\mathbf{c}\sim p(\mathbf{c}),\mathbf{a}\sim\mathcal{N}(0,\mathbf{I}_{a})}\big{[}\big{\lVert}E^{a}\big{(}G^{i}(\mathbf{c,a})\big{)}-\mathbf{a}\big{\rVert}_{1}\big{]}}_{\text{reconstruction of sensitive factors $\mathbf{a}$}}$
$+\underbrace{\mathbb{E}_{\mathbf{m}\sim p(\mathbf{m}),\mathbf{s}\sim\mathcal{N}(0,\mathbf{I}_{s})}\big{[}\big{\lVert}E^{s}(G^{o}(\mathbf{m,s}))-\mathbf{s}\big{\rVert}_{1}\big{]}+\mathbb{E}_{\mathbf{c}\sim p(\mathbf{c}),\mathbf{s}\sim\mathcal{N}(0,\mathbf{I}_{s}),\mathbf{a}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{a})}\big{[}\big{\lVert}E^{s}\big{(}G^{o}(G^{i}(\mathbf{c,a}),\mathbf{s})\big{)}-\mathbf{s}\big{\rVert}_{1}\big{]}}_{\text{reconstruction ofstyle factors $\mathbf{s}$}}$
$+\underbrace{\mathbb{E}_{\mathbf{m}\sim p(\mathbf{m}),\mathbf{s}\sim\mathcal{N}(0,\mathbf{I}_{s})}\big{[}\big{\lVert}E^{m}\big{(}G^{o}(\mathbf{m,s})\big{)}-\mathbf{m}\big{\rVert}_{1}\big{]}}_{\text{reconstruction of semantic factors $\mathbf{m}$ (outer level)}}$
where $p(\mathbf{c})$ is given by $\mathbf{c}=E^{c}(E^{m}(\mathbf{x}))$, $\mathbf{a}=E^{a}(E^{m}(\mathbf{x}))$, and $\mathbf{s}=E^{s}(\mathbf{x})$.
Sensitive Loss. Since a sensitive factor is causally dependent on the sensitive features of a datapoint, as shown in the inner level of Fig. 2, a simple classifier $h:\mathcal{A}\times\Theta\rightarrow\mathcal{Z}$ is trained and further it is used to predict the sensitive label using $\mathbf{a}$ in the second stage.
$\mathcal{L}_{sens}=CrossEntropy(\mathbf{z},\hat{\mathbf{z}})\quad\text{where}\quad\hat{z}=h(\mathbf{a},\bm{\theta}_{z})=h(E^{a}(E^{m}(\mathbf{x})),\bm{\theta}_{z})$

### v2

@@ 6.1 Settings
FEDORA 0.87$\pm$0.09 / 0.53$\pm$0.01 / 62.56$\pm$2.25 0.94$\pm$0.05 / 0.52$\pm$0.01 / 93.36$\pm$1.70 0.93$\pm$0.03 / 0.53$\pm$0.02 / 93.43$\pm$0.73 0.92 / 0.53 / 83.12
6.2 Results
Data augmentation in synthetic domains via $T$. We visualize the augmented samples with random variations in Fig. 4. The first column (Original) shows the images sampled from the datasets. In the second column (Reconstruction), we display images generated from latent factors encoded from the images in the first column. The images in the second column closely resemble those in the first column. Images in the last three columns are generated using the semantic factors encoded from images in the first column, associated with style and sensitive factors randomly sampled from their respective Gaussian distributions. The images in the last three columns preserve the fundamental semantic information of the corresponding samples in the first column. However, their style and sensitive attributes undergo significant changes at random. The generated images within synthetic domains enhance the classifier’s generalization ($f$) to unseen source domains. This demonstrates that the transformation model $T$ effectively extracts latent factors and produces diverse transformations of the provided data domains.
Effectiveness of $T$. To further validate the effectiveness of $T$, drawing inspiration from [48], we train a separate transformation model for each domain. Subsequently, we generate an output image by utilizing distinct latent factors from each domain. Using ccMNIST as an example, we individually train three transformation models $\{T^{i}\}_{i=1}^{3}$ within each domain. Each $T^{i}$ includes unique encoders $E_{c}^{i}$, $E_{a}^{i}$, and $E_{s}^{i}$. As shown in Fig. 4, an output image is generated through $G$ using a semantic factor (digit class, $E_{c}^{1}(\mathbf{x}^{1})$), a sensitive factor (background color, $E_{a}^{2}(\mathbf{x}^{2})$), and a style factor (digit color, $E_{s}^{3}(\mathbf{x}^{3})$) from images in different domains. As a result, the output image is constructed from the digit of $\mathbf{x}^{1}$, the background color of $\mathbf{x}^{2}$, and the digit color of $\mathbf{x}^{3}$, with given variations. This suggests that the augmented data with random variations in Fig. 4 for the synthetic domain are not merely altering colors; instead, they are precisely generated with unchanged semantics and random sensitive and style factors.
The effectiveness of FEDORA across domains in terms of predicted fairness and accuracy. Comprehensive experiments showcase that FEDORA consistently outperforms baselines by a considerable margin. For all tables in the main paper and Appendix, results shown in each column represent performance on the target domain, using the rest as source domains. Due to space limit, selected results for three domains of FairFace are shown in Tab. 4, but the average results are based on all domains. Complete performance for all domains of datasets refers to Appendix E. As shown in Tab. 4, for the FairFace dataset, our method has the best accuracy and fairness level for the average DG performance over all the domains. More specifically, our method has better fairness metrics (3% for $DP$, 2% for $AUC_{fair}$) and comparable accuracy (0.19% better) than the best of the baselines for individual metrics. As shown in Tab. 5, for YFCC100M-FDG, our method excels in fairness metrics (8% for $DP$, 4% for $AUC_{fair}$) and comparable accuracy (0.35% better) compared to the best baselines.
Ablation studies. We conduct three ablation studies to study the robustness of FEDORA on FairFace. In-depth descriptions and the pseudocodes for these studies can be found in Appendix C. More results can be found in Appendix E. (1) In w/o $E_{a}$, we modify the encoder within $T$ by restricting its output to only latent semantic and style factors. (2) w/o $T$ skips data augmentation in synthetic domains via $T$ and results are conducted only based $f$ constrained by fair notions outlined in Defn. 1. (3) In w/o $\mathcal{L}_{fair}$, the fair constraint on $f$ is not included, and we eliminate the $\mathcal{L}_{fair}$ in line 9 of Algorithm 1. We include the performance of such ablation studies in Fig. 6. The results illustrate that when data is disentangled into three factors, and the model is designed accordingly, it can enhance generalization performance due to covariate and dependence shifts. Generating data in synthetic domains with random fairness dependence patterns proves to be an effective approach for ensuring fairness invariance across domains.
Fairness-accuracy tradeoff. In our Algorithm 1, because $\lambda_{2}$ (lines 10 and 12) is the parameter that regularizes the fair loss, we conduct additional experiments to show the change of tradeoffs between accuracy and fairness sweeping over a range of $\lambda_{2}\in[0.01,0.05,0.1,1,10]$. Our results show that the larger (small) $\lambda_{2}$, the better(worse) model fairness for each domain as well as in average, but it gives worse (better) model accuracy. Evaluation on FairFace and YFCC100M-FDG is given in Fig. 6. Results in the top-right of the figure indicate good performance. This result is plotted on the average performance over all target domains.
@@ 7 Conclusion
In this paper, we introduce a novel approach designed to tackle the challenges of domain generalization when confronted with covariate shift and dependence shift simultaneously. We present a tractable algorithm and showcase its effectiveness through comprehensive analyses and exhaustive empirical studies.
@@
Appendix A Notations
@@
$\mathcal{C}$ latent space for semantic factors
$\mathcal{S}$ latent space for style factors
$\mathcal{A}$ latent space for sensitive factors
$\mathbf{c}$ semantic factor
@@
$d[\cdot]$ distance metric over outputs
$dist[\cdot]$ distance metric over distributions
$\mathcal{D}$ data set
@@
$\mathcal{F}$ classifier space
@@
$g(,)$ fairness function
@@
$e$ domain labels
$s$ source domain labels
$\mathcal{E}$ set of data labels
@@
$\delta$ expectation of the relaxed constraint
$h$ sensitive classifier
$\hat{z}$ sensitive attributes predicted by $h$
@@
Appendix B Experimental Settings
B.1 Evaluation Metrics.
Two fair metrics are used for evaluation.
@@ Appendix B Experimental Settings
    $\text{DP}=k,\>\text{if DP}\leq 1;\text{DP}=1/k,\>\text{otherwise}$
    where $k=\mathbb{P}(\hat{Y}=1|Z=-1)/\mathbb{P}(\hat{Y}=1|Z=1)$ This is also known as a lack of disparate impact [64]. A value closer to 1 indicates fairness.
  • The Area Under the ROC Curve ($AUC_{fair}$) [62] varies from zero to one, and it is symmetric around 0.5, which represents random predictability or zero bias effect on predictions.
    $\frac{\sum_{(\mathbf{x}_{i},z=-1,y_{i})\in\mathcal{D}_{-1}}\sum_{(\mathbf{x}_{j},z=1,y_{j})\in\mathcal{D}_{1}}I\big{(}\mathbb{P}(\hat{y}_{i}=1)>\mathbb{P}(\hat{y}_{j}=1)\big{)}}{|\mathcal{D}_{-1}|\times|\mathcal{D}_{1}|}$
@@
B.2 Details of Learning the Transformation Model
For simplicity, we denote the transformation model $T$ consisting of three encoders $E_{c},E_{a},E_{s}$, and a decoder $G$. However, in practice, we consider a bi-level auto-encoder (see Fig. 7), wherein an additional content encoder $E_{m}:\mathcal{X}\rightarrow\mathcal{M}$ takes data as input and outputs a content factor. Furthermore, the decoder $G$ used in the main paper is renamed $G_{o}$. Specifically, the inner level decoder is denoted as $G_{i}:\mathcal{C}\times\mathcal{A}\rightarrow\mathcal{M}$. As a consequence, the transformation model $T$ consists of encoders $E=\{E_{m},E_{s},E_{c},E_{a}\}$ and decoders $G=\{G_{i},G_{o}\}$.
Specifically, in the outer level, an instance is first encoded to a content factor $\mathbf{m}\in\mathcal{M}$ and a style factor $\mathbf{s}\in\mathcal{S}$ through the corresponding encoders $E_{m}$ and $E_{s}$, respectively. In the inner level, the content factor $\mathbf{m}$ is further encoded to a content factor $\mathbf{c}\in\mathcal{C}$ and a sensitive factor $\mathbf{a}\in\mathcal{A}$, through encoders $E_{c}$ and $E_{a}$. Therefore, the bidirectional reconstruction loss and the sensitiveness loss stated in Sec. 4 are reformulated.
$\mathcal{L}_{recon}^{data}=\mathbb{E}_{\mathbf{x}^{s}\sim\mathbb{P}^{s}_{X}}\big{[}\big{\lVert}G_{o}\big{(}\hat{\mathbf{m}},E_{s}(\mathbf{x}^{s})\big{)}-\mathbf{x}^{s}\big{\rVert}_{1}\big{]}+\mathbb{E}_{\mathbf{m}\sim\mathbb{P}_{M}}\big{[}\big{\lVert}G_{i}\big{(}E_{c}(\mathbf{m}),E_{a}(\mathbf{m})\big{)}-\mathbf{m}\big{\rVert}_{1}\big{]}$
where $\hat{\mathbf{m}}=G_{i}(\mathbf{c,a})=G_{i}\big{(}E_{c}(E_{m}(\mathbf{x}^{s})),E_{a}(E_{m}(\mathbf{x}^{s}))\big{)}$; $\mathbb{P}_{M}$ is given by $\mathbf{m}=E_{m}(\mathbf{x}^{s})$.
$\mathcal{L}_{recon}^{factor}=$ $\mathbb{E}_{\mathbf{c}\sim\mathbb{P}_{C},\mathbf{a}\sim\mathcal{N}(0,\mathbf{I}_{a})}\big{[}\big{\lVert}E_{c}\big{(}G_{i}(\mathbf{c,a})\big{)}-\mathbf{c}\big{\rVert}_{1}\big{]}+\mathbb{E}_{\mathbf{c}\sim\mathbb{P}_{C},\mathbf{a}\sim\mathcal{N}(0,\mathbf{I}_{a})}\big{[}\big{\lVert}E_{a}\big{(}G_{i}(\mathbf{c,a})\big{)}-\mathbf{a}\big{\rVert}_{1}\big{]}$
$+\mathbb{E}_{\mathbf{m}\sim\mathbb{P}_{M},\mathbf{s}^{s}\sim\mathcal{N}(0,\mathbf{I}_{s})}\big{[}\big{\lVert}E_{s}(G_{o}(\mathbf{m,s}))-\mathbf{s}\big{\rVert}_{1}\big{]}+\mathbb{E}_{\mathbf{c}\sim\mathbb{P}_{C},\mathbf{s}^{s}\sim\mathcal{N}(0,\mathbf{I}_{s}),\mathbf{a}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{a})}\big{[}\big{\lVert}E_{s}\big{(}G_{o}(G_{i}(\mathbf{c,a}),\mathbf{s})\big{)}-\mathbf{s}\big{\rVert}_{1}\big{]}$
$+\mathbb{E}_{\mathbf{m}\sim\mathbb{P}_{M},\mathbf{s}^{s}\sim\mathcal{N}(0,\mathbf{I}_{s})}\big{[}\big{\lVert}E_{m}\big{(}G_{o}(\mathbf{m,s})\big{)}-\mathbf{m}\big{\rVert}_{1}\big{]}$
where $\mathbb{P}_{C}$ and $\mathbb{P}_{M}$ are given by $\mathbf{c}=E_{c}(E_{m}(\mathbf{x}^{s}))$ and $\mathbf{m}=E_{m}(\mathbf{x}^{s})$. $\mathbf{a}=E_{a}(E_{m}(\mathbf{x}^{s}))$, and $\mathbf{s}=E_{s}(\mathbf{x}^{s})$.
$\mathcal{L}_{sens}=CrossEntropy(z^{s},h(E_{a}(E_{m}(\mathbf{x}^{s}))))$